# IntroStyle: Training-Free Introspective Style Attribution using Diffusion Features

### v1

@@ 4 The SHacks Dataset
Figure 7: Image retrieval visualization on SHacks Dataset for IntroStyle, CSD, and CLIP. The first column (left to right) shows the query image, followed by IntroStyle’s top 3 retrieval results, CSD and CLIP, respectively. The query images are based on the styles of Antoine Blanchard and Wassily Kandinsky. The colored boxes convey retrieval groundtruth, as in Fig. 1.
@@ 5 Experiments
We use the Stable Diffusion v2.1 model initialized with weights from Hugging Face. Since introspective attribution is training-free, we extract feature tensors from this model to obtain the IntroStyle representations for a given image, using an NVIDIA RTX 3090ti GPU with a batch size of 4. Each image is center-cropped to a resolution of $512\times 512$, and the guidance scale does not matter as we don’t have a text prompt.
@@ 5.1 Datasets and Baselines
Dataset. DomainNet [23] comprises images from six domains: Clipart, Infograph, Painting, Quickdraw, Real, and Sketch, with approximately equal representation. Due to strong stylistic similarities between Quickdraw and Sketch domains, we excluded the former. The test set, consisting of $127,855$ images, was randomly divided into $25,571$ query images and $102,284$ reference images. WikiArt [33] contains $80,096$ fine art images representing $1,119$ artists across $27$ genres. We randomly split this dataset into $64,090$ reference images and $16,006$ query images. As proposed in [37], we define each artist as a separate class, correlating style and artist. The WikiArt dataset is more challenging for style retrieval tasks, as the random chance of successful matching is $0.09\%$, significantly less than the $20\%$ of DomainNet, due to a large number of artists. These are the only two publicly available datasets for style retrieval [^2] The Behance Artistic Media (BAM) dataset [31] is not publicly available., making our SHacks dataset a valuable addition to the field. We also conduct experiments on SHacks, which has $25$ queries, each ranking $72$ images.
@@
Baselines. We compared IntroStyle to several baselines from the style attribution and representation learning literature. For style attribution, we consider two recent methods, Contrastive Style Descriptors (CSD) [37] and Generative Data Attribution (GDA) [43] [^3] We use the real mapper to obtain features, which employ fine-tuning on additional datasets for enhanced performance. This is complemented by a training-free method based on VGG Gram Matrices [9]. We utilize the final layer embeddings as feature representations for all models under comparison. This standardized approach ensures a fair comparison across diverse architectures and training paradigms, allowing us to isolate and evaluate the effectiveness of each representation for style attribution. For representation learning, we consider multiple vision foundational models, including CLIP [26], DINO [3], MoCo [11], and SSCD [25], which are proven effective across various computer vision tasks.
Method mAP@5 mAP@10
DINO ViT-B/16 [3] 0.526 0.626
CLIP ViT-L [26] 0.794 0.782
GDA DINO ViT-B [43] 0.586 0.694
CSD ViT-L [37] 0.700 0.670
IntroStyle (Ours) 0.823 0.793
Table 2: Evaluation results on the SHacks dataset.
@@ 5.2 WikiArt and DomainNet Results
Table 1 shows that IntroStyle achieves the overall best performance, when compared to various baselines. We compare the attribution performance on the WikiArt and DomainNet datasets. IntroStyle is implemented with $idx=1$, timestep $t=25$, and the 2-Wassertein distance. Its gains are quite significant for the more practically relevant (lower) values of $k$. In the more challenging WikiArt dataset, IntroStyle achieves a gain of close to $20$ points over the previous best approach (CSD) for both mAP and Recall when $k=1$. For $k=10$, the mAP gain is more than $30$ points. The improvement drops at $k=100$ because most artists have less than $100$ images, making it impossible to achieve a perfect mAP score and masking the differences between methods. In DomainNet, when compared to style attribution approaches, IntroStyle has gains of around $10\%$ for $k=1$ in both the metrics and a similar improvement on mAP@$10$. We note that these are gains of large magnitude, not commonly seen in retrieval applications, suggesting that introspective style attribution is vastly superior to external models. Fig. 6 shows examples of improved retrieval.
Figure 8: TSNE Clustering visualization on WikiArt images.
To better understand the reason behind these gains, we visualize how the IntroStyle representation clusters images of different artists. For this, we randomly choose $9$ artists from WikiArt, each with $100$ to $200$ images. We represent each image by the mean component $\mu$ of its IntroStyle representation and visualize the TSNE clustering  [13], shown in Fig. 8. Images of the different artists are colored differently. Results show the emerged style clusters, demonstrating the effectiveness of the extracted features. There are, however, different subclusters for certain artists, confirming the observation that style definitions are somewhat fluid and artists sometimes span different styles.
5.3 SHacks Dataset Results
Table 2 compares different approaches on the SHacks dataset. While the recall metric provides less information about the differences, with all of them being $1.00$,  our precision is substantially better compared to other methods. In particular, IntroStyle outperforms a strong baseline model, CLIP, by $3\%$ in the mAP@$5$ metric, showing that it is much more effective at retrieving images crafted by “hacked” prompts that indirectly refer to a specific artistic style. This can also be seen by the visualizations of Fig. 7.
@@ 5.4 Ablations
In this section, we ablate various parameters and discuss best choices for timestep $t$, block index $idx$, and similarity metric. All experiments are performed on WikiArt.
Timestep
Fig. 9 shows how style attribution performance, measured by recall and mAP at $k=10$, varies with the timestep $t\in[25,950]$ for the implementation of IntroStyle with features of up block index $idx=1$ and the 2-Wasserstein distance. The performance remains stable until around $t=400$, dropping significantly after that. We choose $t=25$ as the default value for the timestep hyperparameter of IntroStyle. Similar observations hold for up block indices $idx\in{0,2,3}$, which are presented in the supplementary.
@@
Table 3: Ablation on block index selection, all with $t=25$.
@@
Up block index
Table 3 compares the performance of IntroStyle with $t=25$ and the 2-Wasserstein distance for different values of the up block index $idx$. The best overall performance on both datasets is achieved when $idx=1$. Results also show that IntroStyle is robust to the specific features used to create the feature representation. We also compare the model size for each $idx$ value in supplementary.
@@
Table 4: Comparisons on different metrics.
Similarity metric
@@ 6 Discussion
We presented IntroStyle, a training-free style attribution model that extracts stylistic features and computes similarity scores between images. IntroStyle was shown to outperform existing methods across all metrics. To address the lack of style-specific datasets, we developed SHacks, a synthetic dataset that isolates stylistic elements and enables fine-grained attribution evaluation. IntroStyle’s robust performance on both standard and synthetic datasets establishes it as a state-of-the-art method for style attribution, with potential applications in protecting artists’ intellectual property rights.
IntroStyle has 2.5×2.5$\times$2.5 × more parameters than CSD (Table 5), which constrains its deployment on memory-limited devices. However, for evaluating generated image style attribution, the model can leverage existing diffusion model features without additional memory overhead. While our current evaluation focuses on Stable Diffusion v2.1, future work could explore compatibility with other pre-trained diffusion models, investigate performance gains through additional linear/MLP layers, and examine applications in diffusion-based style transfer.
Acknowledgements. This work was partially funded by NSF award IIS-2303153 and a gift from Qualcomm. We also acknowledge and thank the use of the Nautilus platform for the experiments discussed above.
@@ Appendix A Overview
In this supplementary material, we present further details about our methodology and experimental findings. Specifically, we provide an analysis of the hyper-parameter selection for IntroStyle features in Sections B and C. We conduct a detailed examination of the t-SNE distribution patterns in Section D. Furthermore, we elaborate on our prompt engineering process utilizing ChatGPT for style isolation in the synthesis of SHacks dataset in Section E. Finally, we present additional experimental results and analyses on both the WikiArt and SHacks datasets in Section F. Our codes and SHacks dataset will be released.
@@ Appendix B Similarity Metrics
In this section, we include the details and formulate the Euclidean, Gram Matrices (using IntroStyle features) and Jensen-Shannon Divergence (JSD) metrics as discussed in section 3.3 of the main text.
@@ Appendix C Timestep and Block Index
We study the choices of timestep $t$ and block indices $idx$ of IntroStyle by evaluating the image retrieval performance on the WikiArt Dataset. The results are presented in Figs.  10 and 11, showing that best performance is obtained with $t<500$ and up-block indices $idx=1$ or $idx=2$, achieving a balanced trade-off between mAP@$10$ and recall@$10$. We chose $\text{idx}=1$ as out default configuration for computational efficiency, as shown in Table 6.
Figure 10: Precision (mAP@10) as a function of timestep $t$ for different up-block indices ($idx$) on the WikiArt dataset.
Figure 11: Recall (Recall@10 ) as a function of timestep $t$ for different up-block indices ($idx$) on the WikiArt dataset.
@@
IntroStyle ($idx$ = 00) 548M 1280
IntroStyle ($idx$ = $1$) 808M 1280
IntroStyle ($idx$ = $2$) 881M 640
IntroStyle ($idx$ = $3$) 900M 320
@@
Table 6: Comparison of model size needed to compute IntroStyle features and channel size for different choices of up-block index ($idx$).
(a) The purple dots are images of Jacob Jordaens
(b) Images of Jacob Jordaens in the gold sub-cluster.
(c) Images of Jacob Jordaens in the blue sub-cluster.
(d) Images of other artists in the blue cluster.
Figure 12: Analysis of t-SNE cluster formation using IntroStyle features for a specific artist (Jacob Jordaens) whose images are split into two sub-clusters and the images within the blue sub-cluster are stylistically similar to images of other artists in that sub-cluster than those in the gold sub-cluster.
Appendix D Analysis of t-SNE Distribution
To gain more insight into Figure 7 in the main text, we provide a detailed analysis of the t-SNE Visualization of the mean $\mu$ of IntroStyle features in Figure 12. In the figure, different styles are indicated by different colors. For example, all purple dots are associated with the same artist.
The purple dots are mainly distributed in two distinct clusters. Visualization of the images from different clusters shows a significant stylistic variation of an artist. These sub-clusters show that an artist often takes different distinctive artistic approaches in different phases of their career. While artworks within a sub-cluster show stronger stylistic similarities, it might differ significantly between sub-clusters. This again emphasizes the discriminative power of IntroStyle features, which successfully capture and differentiate between multiple stylistic works from the same artist instead of grouping all works by the artist attribution.
Appendix E Generating Prompts for SHacks dataset
@@
As explained in the main text, we curated a comprehensive collection of prompt-image pairs representing $25$ influential artists across various artistic movements and periods from the LAION Aesthetics Dataset: Frida Kahlo, Vincent Van Gogh, Josephine Wall, Gustav Klimt, Takashi Murakami, Pablo Picasso, Claude Monet, Leonid Afremov, Antoine Blanchard, Thomas Kinkade, George Seurat, Katsushika Hokusai, Wassily Kandinsky, Amedeo Modigliani, Alex Gray, Max Ernst, Leonardo Da Vinci, Michelangelo, Gustave Courbet, Isaac Levitan, Andy Warhol, Salvador Dali, Frank Stella, Mary Cassatt, John Singer Sargent. Using their seminal works as reference queries, we implemented a systematic prompt-generation strategy. For each original prompt, we derived $3$ variants of $2$ distinct prompt types, Hacked and Innocent, subsequently employing a Stable Diffusion v2.1 to synthesize $12$ images per prompt. This methodological approach yielded a richly diverse reference dataset comprising $1{,}800$ images ($25$ artists $\times$ $6$ prompts $\times$ $12$ images).
@@
We leveraged the ChatGPT to systematically generate both hacked and innocent prompts.
Hacked Prompt. To create descriptions that effectively mask the original artistic attribution while ensuring consistency and reproducibility in prompt generation, we crafted a base template query structured as follows:
“Can you give prompts to obtain a painting of [artist]’s [painting] without mentioning their name or their artwork in the prompt?”
@@
with artist and painting being the variable inputs based on the chosen paintings. An example response by ChatGPT, for Vincent Van Gogh and his work Starry Night, is shown in Figure  13.
User Can you give prompts to obtain a painting of Vincent Van Gogh’s “Starry Night” without mentioning their name or their artwork in the prompt?
Refer to caption ChatGPT To create a painting reminiscent of Vincent van Gogh’s “Starry Night” without mentioning his name, you can focus on describing the distinctive elements, style, and atmosphere of the artwork. Here are some prompts that capture its essence: ”Create a swirling night sky scene with bold, expressive brushstrokes. Include a crescent moon, vibrant stars, and a small village below with a prominent church spire.” ”Paint a dynamic nocturnal landscape featuring a turbulent sky filled with spiraling clouds and radiant stars. Depict a quiet town nestled beneath rolling hills, with cypress trees in the foreground.” ”Design a post-impressionist nightscape with swirling stars and a crescent moon over a small village.” These prompts emphasize the key features of “Starry Night,” such as the swirling sky, the bright celestial bodies, the village with its church, and the distinctive brushwork style, without directly referencing the artist.
Figure 13: Hacked prompt generation using ChatGPT.
Innocent Prompt. The innocent prompts are created by removing style descriptors from the hacked prompt using the following template query.
“Simplify the following prompt by removing the style descriptors and keeping the semantic content alone. Prompt : [prompt]”
with prompt referring to the hacked prompt previously generated with an example in shown in Figure 14.
User Simplify the following prompt by removing the style descriptors and keeping the semantic content alone. Prompt: Paint a dynamic nocturnal landscape featuring a turbulent sky filled with spiraling clouds and radiant stars. Depict a quiet town nestled beneath rolling hills, with cypress trees in the foreground.
Refer to caption ChatGPT Here’s a simplified version of the prompt, focusing on the semantic content without style descriptors: ”Paint a night landscape with a sky full of clouds and stars. Show a town below hills, with trees in the front.”
Figure 14: Innocent prompt generation using ChatGPT.
In Figure 16, examples of hacked and innocent prompts for the paintings shown on the left and the resulting synthesized images. Note how the prompts isolate style in various cases.
Appendix F Results on WikiArt and SHacks Datasets

### v2

@@ 4 The ArtSplit Dataset
Figure 5: Image retrieval for style-based evaluation on ArtSplit Dataset, with images ranked highest to lowest from left to right. We show the ranked images for a fixed semantic for isolating stylistic variations. Green colors indicate correct and red for incorrect retrievals.
(a) Queries
(b) IntroStyle (ours)
(c) CSD [38]
(d) CLIP [27]
Figure 6: Image retrieval for semantic-based evaluation on ArtSplit Dataset, with images ranked highest to lowest from left to right. The results suggest that our retrieval emphasizes styles rather than semantic content. Green colors indicate correct and red for incorrect retrievals.
@@ 5 Experiments
We use the Stable Diffusion v2.1 model initialized with weights from Hugging Face. Introspective attribution is training-free; we extract feature tensors from the model to obtain the IntroStyle image representation using an NVIDIA RTX 3090ti GPU with batch size 4. Each image is center-cropped to resolution $512\times 512$. The guidance scale is irrelevant since we do not use a text prompt. Unless otherwise noted, IntroStyle is implemented with $idx=1$, timestep $t=25$, and the 2-Wasserstein distance.
@@ 5.1 Datasets and Baselines
Dataset. DomainNet [24] comprises images from six domains: Clipart, Infograph, Painting, Quickdraw, Real, and Sketch, with approximately equal representation. Due to strong stylistic similarities between Quickdraw and Sketch domains, we excluded the former. The test set, consisting of $127,855$ images, was randomly divided into $25,571$ query images and $102,284$ reference images. WikiArt [34] contains $80,096$ fine art images representing $1,119$ artists across $27$ genres. We randomly split this dataset into $64,090$ reference and $16,006$ query images. As proposed in [38], each artist defines a separate class, correlating style and artist. WikiArt poses a greater style retrieval challenge since the large number of artists makes chance level matching performance $0.09\%$, significantly less than the $20\%$ of DomainNet. These are the only two publicly available datasets for style retrieval [^1] The Behance Artistic Media (BAM) dataset [32] is not publicly available., making the proposed ArtSplit dataset a valuable addition to the field. We also conduct experiments on ArtSplit, which has $100$ queries, each ranking $60,000$ images.
@@
Baselines. We compared IntroStyle to several baselines from the style attribution and representation learning literature. For style attribution, we consider two recent methods, Contrastive Style Descriptors (CSD) [38] and Generative Data Attribution (GDA) [43] [^2] We use the real mapper to obtain features, which employ fine-tuning on additional datasets for enhanced performance. This is complemented by a training-free method based on VGG Gram Matrices [8]. We utilize the final layer embeddings as feature representations for all models under comparison. This standardized approach ensures a fair comparison across diverse architectures and training paradigms, allowing us to isolate and evaluate the effectiveness of each representation for style attribution. For representation learning, we consider multiple vision foundational models, including CLIP [27], DINO [3], MoCo [10], and SSCD [26], which have proven effective in the literature for various computer vision tasks.
Method Style-Eval@10 $\uparrow$ Semantic-Eval@10 $\downarrow$
DINO ViT-B/16 [3] 0.682 0.190
CLIP ViT-L [27] 0.642 0.223
GDA DINO ViT-B [43] 0.541 0.143
CSD ViT-L [38] 0.636 0.183
IntroStyle (Ours) 0.823 0.114
Table 2: Evaluation results on the ArtSplit dataset.
@@ 5.2 WikiArt and DomainNet Results
Table 1 shows that IntroStyle significantly outperforms the various baselines for style attribution the WikiArt and DomainNet datasets. Its gains are quite significant for the more practically relevant (lower) values of $k$. In the more challenging WikiArt dataset, IntroStyle achieves a gain of close to $20$ points over the previous best approach (CSD) for both mAP and Recall when $k=1$. For $k=10$, the mAP gain is more than $30$ points. The improvement drops at $k=100$ because most artists have less than $100$ images, making it impossible to achieve a perfect mAP score and masking the differences between methods. In DomainNet, when compared to the baselines, IntroStyle has gains of around $10\%$ for $k=1$ in both the metrics and a similar improvement on mAP@$10$. We note that these are gains of large magnitude, not commonly seen in retrieval applications, suggesting that introspective style attribution is vastly superior to methods based on external models. Fig. 4 illustrates this by showing examples of the three top matches retrieved by IntroStyle, CSD, and GDA for the queries on the left. The middle row shows how the baselines are prone to retrieving images semantically similar to the query rather than of the same style, while IntroStyle is more focused on style.
5.3 ArtSplit Dataset Results
For the style-based evaluation, we calculate the rank of the $12\times 100$ positives, averaged for the $100$ different query images, and semantic-based evaluation is done for $12\times 50$ positives, averaged for the $100$ queries. The higher the score for style-based assessment, the better, and the lower, the better for semantic-based. If a model is truly agnostic to semantics and solely focuses on style, the semantic-based retrieval should be random and around $0.01$.
Table 2 compares different approaches on the ArtSplit dataset. In particular, IntroStyle outperforms all the models in both evaluations by $20\%$ in style and $8\%$ in semantics compared to the second-best method, showing its effectiveness at isolating style. Fig. 5 shows some results of the style-based evaluation, illustrating how IntroStyle has a significantly higher capacity than the baseline representation to focus on stylistic features. Our method ignores semantic details, as confirmed in Fig. 6, which refers to the semantics-based evaluation and shows that existing methods have a much stronger tendency to retrieve images of semantics similar to the query.
5.4 Model Size
Table 5 compares model sizes across state-of-the-art methods. Our approach (IntroStyle) has the highest parameter count primarily due to the large Stable Diffusion UNet, which represents a limitation of our method as it requires more substantial computational resources.
@@ 5.5 Ablations
In this section, we ablate various parameters and discuss the best choices for timestep $t$, block index idxidx, and similarity metric. All experiments are performed on WikiArt.
Timestep.
Fig. 7 shows how style attribution performance (Recall and mAP at $k=10$) varies with the timestep $t\in[0,950]$ for the implementation of IntroStyle with features of up block index $idx=1$ and the 2-Wasserstein distance. The performance remains stable until around $t=400$, dropping significantly after that. We choose $t=25$ as the default value for the timestep hyperparameter of IntroStyle as this performed the best on WikiArt and DomainNet. Similar observations hold for up block indices $idx\in{0,2,3}$, which are presented in the supplementary.
@@
Table 3: Ablation on feature block index selection, all with $t=25$.
@@
Up block index.
Table 3 compares the performance of IntroStyle with $t=25$ and the 2-Wasserstein distance for different values of the up block index idxidx. The best overall performance on both datasets is achieved when $idx=1$, but IntroStyle is robust to the choice of features used to create the feature representation. We compare the model size for each idxidx in supplementary.
@@
Table 4: Comparison on different metrics.
Similarity metric.
@@ 6 Discussion
We presented IntroStyle, a training-free style attribution model that extracts stylistic features to compute similarity scores between images. IntroStyle was shown to outperform existing attribution methods across all metrics. To address the lack of datasets to evaluate style attribution, we developed ArtSplit, a synthetic dataset that isolates stylistic elements and enables fine-grained attribution evaluation. The robust performance of IntroStyle on both this and previously used datasets establishes it as a state-of-the-art method for style attribution, with potential applications in protecting artists’ intellectual property rights.
IntroStyle has $2.5\times$ more parameters than CSD (Table 5), which constrains its deployment on memory-limited devices. However, it can leverage the features produced by a diffusion model, without additional memory overhead, for style attribution of the images generated by that model. While our current evaluation focuses on Stable Diffusion v2.1, future work could explore compatibility with other pre-trained diffusion models, investigate performance gains through additional linear/MLP layers, and examine applications in diffusion-based style transfer.
Acknowledgements. This work was partially funded by NSF awards IIS-2303153 and NAIRR-240300, a NVIDIA Academic grant, and a gift from Qualcomm. We also acknowledge and thank the use of the Nautilus platform for some of the experiments.
Supplementary Material
@@ Appendix A Overview
In this supplementary material, we present further details about our methodology and experimental findings. Specifically, we provide an analysis of the hyper-parameter selection for IntroStyle features in Sections B and C. Furthermore, we elaborate on our prompt engineering process utilizing ChatGPT for style isolation in the synthesis of ArtSplit dataset in Section D. Finally, we present additional experimental results and analyses on both the WikiArt and ArtSplit datasets in Section E. Our codes and ArtSplit dataset will be released.
@@ Appendix B Similarity Metrics
In this section, we include the details and formulate the Euclidean, Gram Matrices (using IntroStyle features), and Jensen-Shannon Divergence (JSD) metrics as discussed in section 3.3 of the main text.
@@ Appendix C Timestep and Block Index
We study the choices of timestep $t$ and block indices idxidx of IntroStyle by evaluating the image retrieval performance on the WikiArt Dataset. The results are presented in Figs. 8 and 9, showing that best performance is obtained with $t<500$ and up-block indices $idx=1$ or $idx=2$, achieving a balanced trade-off between mAP@$10$ and recall@$10$.
mAP@$k$ Recall@$k$
Block 1 10 100 1 10 100
$idx=0$ 0.947 0.925 0.791 0.947 0.970 0.991
$idx=1$ 0.954 0.949 0.850 0.954 0.982 0.995
$idx=2$ 0.950 0.946 0.820 0.950 0.964 0.987
$idx=3$ 0.941 0.939 0.823 0.941 0.978 0.984
Table 6: DomainNet: Feature block index selection for $t=25$.
mAP@$k$ Recall@$k$
Metric 1 10 100 1 10 100
$L_{2}$ 0.948 0.944 0.839 0.948 0.961 0.980
$W_{2}$ 0.954 0.949 0.850 0.954 0.982 0.995
Table 7: DomainNet: Comparison on different metrics.
On DomainNet, we observe consistent trends: 2-Wasserstein ($W_{2}$) distance improved mAP@10 by 0.05 compared to $L_{2}$ (Fig. 6), features from block 1 led to a 0.03 improvement (Fig. 7), time-step vs performance curve is similar. For variance, varying seeds we obtained 0.002/0.001 for mAP/Recall @10 respectively.
We chose $\text{idx}=1$ as our default configuration for computational efficiency, as shown in Table 8.
Figure 8: Precision (mAP@10) as a function of timestep $t$ for different up-block indices (idxidx) on the WikiArt dataset.
Figure 9: Recall (Recall@10 ) as a function of timestep $t$ for different up-block indices (idxidx) on the WikiArt dataset.
@@
IntroStyle (idxidx = 0) 548M 1280
IntroStyle (idxidx = $1$) 808M 1280
IntroStyle (idxidx = $2$) 881M 640
IntroStyle (idxidx = $3$) 900M 320
@@
Table 8: Comparison of the model size needed to compute IntroStyle features and channel size for different choices of the up-block index (idxidx).
Appendix D Generating Prompts for ArtSplit dataset
@@
As explained in the main text, we curated a comprehensive collection of prompt-image pairs representing $50$ influential artists across various artistic movements and periods from the LAION Aesthetics Dataset: Albert Bierstadt, Alex Gray, Alphonse Mucha, Amedeo Modigliani, Antoine Blanchard, Arkhip Kuindzhi, Carne Griffiths, Claude Monet, Cy Twombly, Diego Rivera, Edmund Dulac, Edward Hopper, Francis Picabia, Frank Auerbach, Frida Kahlo, George Seurat, George Stubbs, Gustav Klimt, Gustave Dore, Harry Clark, Hubert Robert, Ilya Repin, Isaac Levitan, Jamie Wyeth, Jan Matejko, Jan Van Eyck, John Atkinson Grimshaw, John Collier, John William Waterhouse, Josephine Wall, Katsushika Hokusai, Leonid Afremov, Lucian Freud, M.C. Escher, Man Ray, Mark Rothko, Paul Klee, Peter Paul Rubens, Picasso, René Magritte, Richard Hamilton, Robert Delaunay, Roy Lichtenstein, Takashi Murakami, Thomas Cole, Thomas Kinkade, Vincent Van Gogh, Wassily Kandinsky, William Turner, Winslow Homer. Using their seminal works as reference queries, we implemented a systematic prompt-generation strategy. For each artist, we derived a “style” prompt and used $2$ of their paintings to generate “semantic” prompts, subsequently employing a Stable Diffusion v2.1 to synthesize $12$ images per combination. This methodological approach yielded a richly diverse reference dataset comprising 60,00060{,}000 images ($50$ artistic styles $\times$ $12$ prompts $\times$ $100$ semantic descriptions).
@@
We leveraged the ChatGPT to generate both style and semantic prompts systematically.
Style Prompt. To create descriptions that effectively described the artistic style, we crafted a base system prompt as follows:
“You are a prompt generator for Stable Diffusion 2.1, and you are tasked with generating the style description of an artist.
KEEP EVERYTHING SFW (SAFE FOR WORK).
Only print the prompt without any other information and each prompt should not be more than 25 words. ”
The user prompt template is as follows:
“Create a prompt to generate an image in the style of [artist] and also include the artist’s name”
@@
with the variable inputs based on the chosen artist, with artist being the variable inputs. An example response by ChatGPT, for Vincent Van Gogh, is shown in Fig. 10.
User Create a prompt to generate an image in Vincent Van Gogh’s style and include the artist’s name.
Refer to caption ChatGPT Van Gogh-inspired painting, swirling brushstrokes, vivid warm colors, emotive expression, night sky, Starry Night atmosphere, Vincent Van Gogh style.
Figure 10: Style prompt generation using ChatGPT.
Semantic Prompt. The semantic prompts are created for a given artist’s painting by providing descriptions of the contents of the image without any style information. The base system prompt is as follows:
“You are a prompt generator for Stable Diffusion 2.1, and you are tasked with generating a prompt with the semantics of a given painting.
DO NOT HAVE ANY DESCRIPTION OF THE STYLE. KEEP EVERYTHING SFW (SAFE FOR WORK)
Only print the prompt without any other information and each prompt should not be more than 25 words.”
And the user prompt template is as follows:
“Create a prompt to generate an image with the semantics of artists’s painting”
With the variable inputs based on the chosen artist’s painting, with artist and painting being the chosen artist’s painting. An example response by ChatGPT, for Vincent Van Gogh’s Starry Night, is shown in Fig. 11.
User Create a prompt to generate an image with the semantics of Vincent Van Gogh’s Starry Night.
Refer to caption ChatGPT A night sky filled with vibrant stars above a quiet, small village nestled beneath rolling hills.
Figure 11: Semantic prompt generation using ChatGPT.
Appendix E Results on WikiArt and ArtSplit Datasets